### v1

# The Emergence of Reproducibility and Consistency in Diffusion Models

Abstract
Recently, diffusion models have emerged as powerful deep generative models, showcasing cutting-edge performance across various applications such as image generation, solving inverse problems, and text-to-image synthesis. These models generate new data (e.g., images) by transforming random noise inputs through a reverse diffusion process. In this work, we uncover a distinct and prevalent phenomenon within diffusion models in contrast to most other generative models, which we refer to as “consistent model reproducibility”. To elaborate, our extensive experiments have consistently shown that when starting with the same initial noise input and sampling with a deterministic solver, diffusion models tend to produce nearly identical output content. This consistency holds true regardless of the choices of model architectures and training procedures. Additionally, our research has unveiled that this exceptional model reproducibility manifests in two distinct training regimes: (i) “memorization regime,” characterized by a significantly overparameterized model which attains reproducibility mainly by memorizing the training data; (ii) “generalization regime,” in which the model is trained on an extensive dataset, and its reproducibility emerges with the model’s generalization capabilities. Our analysis provides theoretical justification for the model reproducibility in “memorization regime”. Moreover, our research reveals that this valuable property generalizes to many variants of diffusion models, including conditional diffusion models, diffusion models for solving inverse problems, and fine-tuned diffusion models. A deeper understanding of this phenomenon has the potential to yield more interpretable and controlable data generative processes based on diffusion models.
1 Introduction
Recently, diffusion models have emerged as a powerful new family of deep generative models with remarkable performance in many applications, including image generation [1, 2, 3] , image-to-image translation [4, 5, 6], text-to-image synthesis [3, 7, 8], and solving inverse problem solving [9, 10, 11, 12]. These models learn an unknown data distribution generated from the Gaussian noise distribution through a process that imitates the non-equilibrium thermodynamic diffusion process [1, 2]. In the forward diffusion process, the noise is continuously injected into training samples; while in the reverse diffusion process, a model is learned to remove the noise from noisy samples parametrized by a noise-predictor neural network. Then guided by the trained model, new samples (e.g., images) from the target data distribution can be generated by transforming random noise instances through step-by-step denoising following the reverse diffusion process. Despite the remarkable data generation capabilities demonstrated by diffusion models, the fundamental mechanisms driving their performance are largely under-explored.
In this work, to better understand diffusion models, we study the following fundamental question: {tcolorbox}
Q1: Starting from the same noise input, how are the generated data samples from various diffusion models related to each other?
A deeper understanding of this question has the potential to yield more interpretable and controlled data generative processes of using the diffusion models in many application disciplines [13, 14]. For example, in the text-driven image generation, insights into this question could help to guide the content generation (e.g., adversial attacking [15], robust defending [16], copyright protection [17, 18]) using the same text embedding but varying noise inputs. In solving inverse problems, an answer to this question will guide us to select the input noise for reducing the uncertainty and variance in our signal reconstruction  [19, 20, 21]. Theoretically, understanding the question will shed light on how the mapping function is learned and constructed between the noise and data distributions, which is crucial for understanding the generation process through the distribution transformation or identifiable encoding [22, 23, 24].
Figure 1: Unconditional diffusion model samples visualization. We utilized denoising diffusion probabilistic models (DDPM) [1, 25], consistency model (CT) [26], U-ViT [27] trained on CIFAR-10 [28] dataset released by the author. Samples in the corresponding row and column are generated from the same initial noise with a deterministic ODE sampler. More details are shown in Table 1 in Appendix A.
In this work, we provide an in-depth study of the question and uncover an intriguing and prevalent phenomenon within the diffusion model that sets it apart from most other generative models. We term this phenomenon as “consistent model reproducibility”. More precisely, as illustrated in Figure 1, when different diffusion models are trained on the same dataset and generate new samples through a deterministic ODE sampler from the same noises, we find that all diffusion models generate nearly identical images, which is irrespective of network architectures, training and sampling procedures, and perturbation kernels.
The consistent model reproducibility we identified for diffusion models is similar to the notion of unique identifiable encoding for deep latent-variable models, which is the property that the learned input-embedding is reproducible towards an identifiable mapping, regardless of different weight initialization or optimization procedures [22]. The property was first proved by [29, 30, 31] on the analysis of Independent Component Analysis (ICA) and extended to deep latent-variable models. Recently, [23] demonstrated the identifiability of Variational Autoencoder (VAE) based on conditionally factorial priors distribution over the latent variables, [24] consider the identifiability theory for conditional energy-based models and [22] proved a linear identifiability on representation learning. In comparison, the consistent model reproducibility of diffusion models studied in this work implies that we are learning a unique encoding between the noise space and the image space.
More interestingly, as illustrated in Figure 2, we find that the consistent model reproducibility of diffusion models emerges in two distinct regimes: (i) "memorization regime" where the model has large capacity of memorizing the training data but no ability to generate new samples, and (ii) "generalization regime" where the model regain the consistent model reproducibility and can also produce new data. In this work, we provide theoretical justification for the memorization regime. Moreover, the simultaneous occurrence of reproducibility and generalizability in the model generalization regime presents an interesting open question that is worth of future study.
(a) Reproducibilty
(b) Generalizability
Figure 2: “Memorization” and “Generalization” regimes for unconditional diffusion models. We utilize DDPMv4 and train them on the CIFAR-10 dataset, adjusting both the model’s size and the size of the training dataset. In terms of model size, we experiment with UNet-64, UNet-128, and UNet-256, where, for instance, UNet-64 indicates a UNet structure with an embedding dimension of 64. As for the dataset size, we select images from the CIFAR dataset, ranging from $2^{6}$ to $2^{15}$. Under each dataset size, different models are trained from the same subset of images. The figure on the left displays the reproducibility score as we compare various models across different dataset sizes, while the figure on the right illustrates the generalizability score of the models as the dataset size changes.
Finally, we show that the consistent model reproducibility could be generalized to many other variants of diffusion model settings: conditional diffusion models, diffusion models for solving inverse problem, and fine-tuning diffusion models. Regarding conditional diffusion models, the model reproducibility is not only evident among different conditional diffusion models, but also it manifests in a structured way when comparing conditional models to their unconditional counterparts. Concerning solving inverse problem by using diffusion models as generative priors, we observe that the model reproducibility is confined within models using the same type of network architectures (e.g., either U-Net based or transformer-based architecture). Finally, in the context of fine-tuning diffusion models, we show partial fine-tuning of pretrained models reduces reproducibility but improves generalizability in “memorization regime” when compared to training from scratch.
Summary of Contributions.
In summary, we briefly highlight our contributions below:
  • Model reproducibility of unconditional diffusion models (Section 2). We provide a systematic study of the of the unconditional diffusion model, regardless of the choices of network architectures, perturbation kernels, training and sampling settings.
  • Two regimes of model reproducibility (Section 3). We find a strong correlation between reproducibility and generalizability for diffusion model. We provide a theoretical study for the memorization regime.
  • Model reproducibility of variants of diffusion models (Section 4). We reveal the reproducibility in more various diffusion model settings, including conditional diffusion models, diffusion models for inverse problem solving, and fine-tuning diffusion models.
Finally, we conclude and discuss the implications of our results in Section 5.
2 Study for Unconditional diffusion models
For unconditional diffusion models, new samples (e.g., images) are generated by transforming randomly sampled noise instances through a reverse diffusion process guided by the trained model [1, 2]. If we start from the same noise input and use a deterministic ODE sampler, we observe that {tcolorbox}
C2: Diffusion models consistently generate nearly identical contents, irrespective of network architectures, training and sampling procedures, and perturbation kernels.
This seems to be obvious by examining the visualized samples generated by different diffusion models in Figure 1 – starting from the same noise input, different diffusion models (i.e., DDPM [1], Consistency Training (CT) [26], and U-ViT [27]) generate nearly identical samples with very similar low-level color structures. More recent seminal [2] has observed a similar phenomenon (see also subsequent works [26, 32]), but the study in [2] remains preliminary.
2.1 Evaluation Setup of Model Reproducibility
While the findings in Figure 1 are intriguing, the basic visualization alone is not sufficient to fully justify our claim. For a more comprehensive study, we first introduce quantitative measures for model reproducibility and basic experimental setup.
2.1.1 Quantitative Measures of Model Reproducibility
To study this phenomenon more quantitatively, we introduce the reproducibility (RP) score to measure the similarity of image pair generated from two different diffusion models starting from the same noise drawn i.i.d. from the standard Gaussian distribution. Specifically, we define
$\text{RP Score}\;:=\;\mathbb{P}\left(\mathcal{M}_{\text{SSCD}}(\bm{x}_{1},\bm{x}_{2})>0.6\right),$
represents the probability of a generated sample pair $(\bm{x}_{1},\bm{x}_{2})$ from two different diffusion models to have self-supervised copy detection (SSCD) similarity $\mathcal{M}_{\text{SSCD}}$ larger than $0.6$ [33, 17] ($\mathcal{M}_{\text{SSCD}}>0.6$ exhibits strong visual similarities). We sampled 10K noise to estimate the probability. The SSCD similarity is first introduced in [33] to measure the replication between image pair $(\bm{x}_{1},\bm{x}_{2})$, which is defined as the following:
$\mathcal{M}_{\text{SSCD}}(\bm{x}_{1},\bm{x}_{2})=\dfrac{\text{SSCD}(\bm{x}_{1})\cdot\text{SSCD}(\bm{x}_{2})}{||\text{SSCD}(\bm{x}_{1})||_{2}\cdot||\text{SSCD}(\bm{x}_{2})||_{2}}$
where $\text{SSCD}(\cdot)$ represents a neural descriptor for copy detection.
In addition, we also use the mean-absolute-error (MAE) score to measure the reproducibility, $\text{MAE Score}:=\mathbb{P}\left(\text{MAE}(\bm{x}_{1},\bm{x}_{2})<15.0\right)$ , similar setting with the RP score. $\text{MAE}(\cdot)$ is the operator that measures the mean absolute different of image pairs in pixel value space ([0, 255]). The quantitative results for selected diffusion model architectures are shown in Figure 3.
2.1.2 Experimental Setup for Evaluation
In this work, we conduct a comprehensive study of model reproducibility based upon different network architectures, model perturbation kernels, and training and sampling processes:
  • Network architectures. We evaluate for both UNet [34] based architecture: DDPM [1], DDPM++ [2], Multistage [35], EDM [32], Consistency Training (CT) and Distillation (CD) [26]; Transformer [36] based architecture: DiT [37] and U-ViT [27].
  • Training Process. We also considered discrete [1] or continuous [2] settings, estimating noise $\epsilon$ or original image $\bm{x}_{0}$, training from scratch or distillation [38, 26] for the diffusion model.
  • Perturbation Kernels. For noise perturbation stochastic differential equations, we use Variance Preserving (VP) [1], Variance Exploding (VE), sub Variance Preserving (sub-VP) [2]. All the models are trained with the CIFAR-10 dataset [28] until convergence. [^2] Here, convergence means achieving the lowest Fréchet inception distance (FID) [39].
  • Sampling Process. For sampling, we only used the deterministic sampler, such as DPM-Solver [40], Heun-Solver [32], DDIM [25] etc. We use the deterministic sampler instead of the stochastic sampler, because randomness introduced by stochastic samplers will break the model reproducibility.
In the process of sampling, it’s important to not that we exclusively use deterministic samplers for two main reasons. First, stochastic samplers (e.g. [1]) introduce random noise at every time step, leading to a lack of model reproducibility. Second, many recent advanced samplers, such as those proposed in [25, 40, 32, 41] are deterministic samplers. These deterministic samplers lead to fast sampling speed, and have been successfully applied not only to unconditional diffusion models but also to diffusion models for solving inverse problems [12], large text-to-image diffusion model [3] and so on.
2.2 Model Reproducibility of Unconditional Diffusion Models
Figure 3: Similarity among different unconditional diffusion model settings. We visualize the quantitative results based upon seven different unconditional diffusion models (DDPMv4, DDPMv6 [1, 25], Multistagev1 [35], EDMv1 [32], UViT [27], CT [26], Progressivev1 [38]) based upon reproducibility score (left) and MAE score (right) (defined in Section 2.1.1). About more detailed settings and a more comprehensive compare could be found in Appendix A.
2.2.1 Quantitative Studies of Model Reproducibility
Based upon the setup in Section 2.1, we provide a quantitative study by comparing the similarity metrics of samples generated based upon different network architectures, model perturbation kernels, and training process. In Figure 3, we visualize the similarity matrix for seven different popular diffusion models, where each element of the matrix measures pairwise similarities of two different diffusion models based upon reproducibility score (left) and MAE score (right). More comprehensive studies are shown in Appendix A. As we can see, there is a very consistent model reproducible phenomenon for comparing any two models. For even the most dissimilar models, (i.e., CT and UViT), the RP and MAE scores are notably high at 0.7 and 0.64, respectively. Specifically, we draw the following conclusions:
  • Model reproducibility is consistent among different network architectures. When comparing DDPMv4, Multistagev1, EDMv1, and UViT, this phenomenon remains consistent regardless of the specific architecture employed.
  • Model reproducibility is consistent among different training procedures. When we compare CT (consistency loss) and EDMv1 (diffusion model loss), even when using different loss functions, they ultimately converge to a similar noise-to-image mapping. Notably, EDMv1 employs a distinct loss by estimating the clear image $\bm{x}$ whereas other methods estimate the noise $\bm{\epsilon}$. Additionally, comparing DDPMv1 and Progressivev1 reveals that both training from scratch and distillation approaches lead to the same results.
  • Model reproducibility is consistent among different sampling procedures. DDPMv4 utilizes DPM-solver, EDMv1 employs a 2nd order heun-solver, and CT utilizes consistency sampling, yet they all exhibit reproducibility.
  • Model reproducibility is consistent among different perturbation kernels. When compared to DDPMv4, DDPMv6, and EDMv1, the reproducibility remains unaffected by the choice of perturbation method (VP, sub-VP, and VE, respectively).
Figure 4: Unqiue Encoding from Noise Hyperplane to Image Manifold. The diagram illustrates the process of mapping from a noise hyperplane to the image manifold. We employ three distinct models: DDPMv6, EDMv1, and Multistagev1. Initially, we select three different initial noises from Gaussion and generate corresponding samples, denoted by a triangle, square, and circle in the first three images on the left. The hyperplane is defined based on these chosen noises. Image generations, starting from uniformly selected initial noise within this hyperplane, are classified as identical to either the triangle, square, or circle image, determined by the maximum SSCD similarity with them. Each initial noise is colored according to its generation’s corresponding class (as indicated on the right; for instance, the noise’s generation identical to the triangle image is represented by the black-white color bar), along with the SSCD similarity to the identical image.
2.2.2 Studies of the Unqiue Encoding from Noise Hyperplane to Image Manifold
The model reproducibility of diffusion models implies that we are learning a unique decoding $f:\mathcal{E}\mapsto\mathcal{I}$ from the gaussian noise space $\mathcal{E}$ to the image manifold $\mathcal{I}$. Specifically, we find that
  • Similar unique encoding maps across different network architectures. We further confirm the model reproducibility by visualizing the mapping $f$ from a 2D noise hyperplane $\mathcal{H}\subseteq\mathcal{E}$ to the image manifold $\mathcal{I}$, inspired by [42]. The visualization in Figure 4 shows that different mappings of different network architectures share very similar structures.
  • Local Lipschitzness of the unique encoding from noise to image space. Furthermore, our visualization suggests that the unique encoding $f$ is locally Lipschitz, where $\|f(\bm{\epsilon}_{1})-f(\bm{\epsilon}_{2})\|\leq L\|\bm{\epsilon}_{1}-\bm{\epsilon}_{2}\|$ for any $\bm{\epsilon}_{1},\bm{\epsilon}_{2}\in\mathcal{B}(\bm{\epsilon},\delta)\cap\mathcal{E}$ with some Lipschitz constant $L$. Here $\mathcal{B}(\bm{\epsilon},\delta)$ denotes a ball centered at a Gaussian noise $\bm{\epsilon}$ with radius $\delta$. In other words, noises $\bm{\epsilon}_{1},\bm{\epsilon}_{2}\in\mathcal{E}$ close in distance would generate similar reproducible images in $\mathcal{I}$ via diffusion models.
Specifically, the visualization in Figure 4 is created as follows. First, we pick three initial noises $(\bm{\epsilon}_{1},\bm{\epsilon}_{2},\bm{\epsilon}_{3})$ in the noise space $\mathcal{E}$ and used different diffusion model architectures to generate clear images ($\bm{x}_{1}$, $\bm{x}_{2}$, $\bm{x}_{3}$) in the image manifold $\mathcal{I}$, so that the images $\{\bm{x}_{i}\}_{i=1}^{3}$ belong to three different classes. Second, we create a 2D noise hyperplane with
$\bm{\epsilon}\left(\alpha,\beta\right)=\alpha\cdot(\bm{\epsilon}_{2}-\bm{\epsilon}_{1})+\beta\cdot(\bm{\epsilon}_{3}-\bm{\epsilon}_{1})+\bm{\epsilon}_{1}$
Within the region $(\alpha,\beta)\in[-0.1,1.1]\times[-0.1,1,1]$, we uniformly sample $100$ points along each axis and generate images $\bm{x}\left(\alpha,\beta\right)$ for each sample $\bm{\epsilon}\left(\alpha,\beta\right)$ using different diffusion model architectures (i.e., DDPMv6, EDMv1, Multistagev1). For each point $\left(\alpha,\beta\right)$, it is considered as identical to image $\bm{x}_{i}$ for $i=\operatorname*{\text{arg~{}max}}_{k\in\{1,2,3\}}[\mathcal{M}_{\text{SSCD}}(\bm{x}_{k},\bm{x}\left(\alpha,\beta\right))]$, and we visualize the value of $\mathcal{M}_{\text{SSCD}}(\bm{x}_{i},\bm{x}\left(\alpha,\beta\right))$. As we observe from Figure 4, the visualization shares very similar structures across different network architectures. Second, for each plot, closeby noises create images with very high similarities. These observations support our above claims.
2.2.3 Comparison with Other Types of Generative Models
We end this section by highlighting that only diffusion models appear to consistently exhibit model reproducibility. This property is seemingly absent in other generative models with one exception as noted in [23].333[23] demonstrates that VAE is uniquely identifiable encoding given a factorized prior distribution over the latent variables. In our evaluation of GAN-based methods, we contrast two variants that have been widely used: Wasserstein GAN (wGAN) [43] and Spectral Normalization GAN (SNGAN) [44]. We conduct this analysis using the CIFAR-10 dataset. In the meanwhile, for VAE-based approaches, we consider both the standard VAE and the Variational Autoencoding Mutual Information Bottleneck (VAMP) method [45]. We evaluate both VAE-based models on the MNIST dataset [46]. The similarity matrices, presented in Figure 5, indicate a lack of model reproducibility in both GAN and VAE based methods.
Figure 5: Quantitative results for GANS and VAEs. We apply SNGAN, WGAN for the reproducibility evaluation of GAN-based approaches and VAE, VAMP for VAE-based approaches. Each model utilized in this analysis was provided by its respective author, and the reproducibility score is defined the same as in previous experiments. Particularly, the latent space for VAE-based methods is learned through the encoder. In this context, our approach involves sampling initial noise from the latent space of VAE and employing it for the generation of VAMP.
3 Correlation between Reproducibility & Generalizability
More interestingly, both of our empirical and theoretical studies in this section demonstrate that {tcolorbox}
C3: The consistent model reproducibility of diffusion models manifests in two distinct training regimes, both strongly correlated with the model’s generalizability.
Here, the model’s generalizability means the model’s ability to generate new samples that is different from samples in the training dataset. Specifically, we measure the generalizability by introducing the generalization (GL) score $:=1-\mathbb{P}\left(\max_{i\in[N]}\left[\mathcal{M}_{\text{SSCD}}(\bm{x},\bm{y}_{i})\right]>0.6\right)$ represents one minus the probability of maximum $\mathcal{M}_{\text{SSCD}}$ is larger than $0.6$, between the generated sample $\bm{x}$ from one diffusion model and all samples $\bm{y}_{i}$ from training dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$ of $N$-samples . We sampled 10K initial noise to estimate the probability.
3.1 Two Regimes of Model Reproducibility
Now, we formally introduce the two regimes of model reproducibility for diffusion models in the following. Both of them exhibit a strong correlation with the model’s ability to generalize.
  • “Memorization regime” characterizes a scenario where the trained model has large model capacity than the size of training data. In this regime, starting from the same noise the model possesses the ability to reproduce the same results, as illustrated in the left region of Figure 1(a) with small training data size. However, the generated samples are often replications of the samples in the training data, and the model lacks the full capacity to generate new samples; see the left region of Figure 1(b). Therefore, we call this regime the “memorization regime” as the generated samples are only replication of training data. In this regime, we can rigorously characterize the optimal denoiser as shown in Theorem 1, and in Figure 6 we empirically demonstrated that the diffusion models converge to the theoretical solutions when the model is highly overparameterized. However, given no generalizability, training diffusion models in this regime holds limited practical interest.
  • “Generalization regime” emerges when the diffusion model is trained on large dataset and doesn’t have full capacity to memorize the whole dataset. Remarkably, as shown in the right region of Figure 1(a), the diffusion model regains the reproducibility as the ratio between dataset size and model capacity increases. Simultaneously, the model possess the ability to generate new samples, as illustrated in the right region of Figure 1(b); this generalization phenomenon is also observed in [47]. As such, we call this regime the “generalization regime” because the stage of generalization is in coincidence with the stage of model reproducibility. This is the regime in which diffusion models are commonly trained and employed in practice, and the coexistence of model reproducibility and generalizability is an intriguing phenomenon to be further understood theoretically.
3.2 A Theoretical Study of the Memorization Regime
For the rest of this section, we provide a theoretical study of the memorization regime, and we leave the study of generalization regime for future work. Specifically, we characterize the optimal denoiser of the diffusion model and show that it results in a unique identifiable encoding under proper assumptions.
Figure 6: Experiment verification of the theory. We employ DDPMv4 and conduct training on the CIFAR-10 dataset. During this process, we make modifications to both the model’s capacity and the size of the training dataset, maintaining the same configuration as depicted in Figure 2. The figure presented here illustrates the reproducibility score between each diffusion model and the theoretically unique identifiable encoding as outlined in Theorem 1.
Theorem 1.
Suppose we train a diffusion model denoiser function $\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)$ with parameter $\bm{\theta}$ on a training dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$ of $N$-samples, by minimizing the training loss
[…]
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}\left[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}\right].$ (2)
Moreover, suppose a trained diffusion model could converge to the optimal denoiser $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ and we use a deterministic ODE sampler to generate images using $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$, then $f:\mathcal{E}\mapsto\mathcal{I}$, which is determined by the $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ and the ODE sampler, is an invertiable mapping and the inverse mapping $f^{-1}$ is a unique identifiable encoding.
The proof for Theorem 1 can be found in the Appendix B, building upon previous findings from [32, 48]. It is worth noting that the optimal denoiser $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ in (2) is deterministic for a given training dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$ and for specific perturbation parameters $s_{t},\sigma_{t}$. Moreover, [48] demonstrates optimal denoiser lacks generalizability, as all the samples generated from $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ are confined within the training dataset. Furthermore, we verified our theory experimentally by comparing the reproducibility score between the the theoretical generations in Theorem 1 and that of a trained network in practice. This can be illustrated by the left region of Figure 6, the trained networks have a very high similarity compared with the theoretical solution on small training dataset, and the similarity decreases when the size of training data increases. As such, we conclude that the diffusion model could converge to the theoretical solution when the model capacity is large enough.
4 Study Beyond Unconditional Diffusion Models
Furthermore, in this section, we explore how model reproducibility can be generalized to variations of the diffusion models, showcasing that: {tcolorbox}
C4: Model reproducibility holds more generally across conditional diffusion models, diffusion models for inverse problems, the fine-tuning of diffusion models.
4.1 Model Reproducibility in Conditional Diffusion Models
Conditional diffusion, introduced by [49, 50], gained its popularity in many applications such as text-to-image generation [3, 7, 8] with superior performance, through the fusion of rich class embeddings with the denoiser function. Interestingly, we find that:
Model reproducibility of conditional models is evident and linked with unconditional counterparts.
Specifically, our experiments in Figure 7 demonstrate that (i) model reproducibility exists among different conditional diffusion models, and (ii) model reproducibility is present between conditional and unconditional diffusion models only if the type (or class) of content generated by the unconditional models matches that of the conditional models. More results can be found in Appendix C.
To support our claims, we introduce the conditional reproducibility score between different conditional diffusion models by $\text{RP}_{cond}\text{ Score}:=\mathbb{P}\left(\mathcal{M}_{\text{SSCD}}(\bm{x}^{c}_{1},\bm{x}^{c}_{2})>0.6\mid c\in\mathcal{C}\right)$, where the pair $(\bm{x}^{c}_{1},\bm{x}^{c}_{2})$ are generated by two conditional models from the same initial noise and conditioned on the class $c\in\mathcal{C}$. Additionally, the between reproducibility score of conditional and unconditional diffusion models is defined as $\text{RP}_{between}\text{ Score}:=\mathbb{P}\left(\max_{{}_{c\in\mathcal{C}}}[\mathcal{M}_{\text{SSCD}}(\bm{x}_{1},\bm{x}^{c}_{2})]>0.6\right)$, for an unconditional generation $\bm{x}_{1}$ and conditional generation $\bm{x}^{c}_{2}$ starting from the same noise. First, when conditioned on the same class and initial noise, results in Figure 7 (a) (b) highlight the similarity of samples generated from different condition diffusion models (i.e., EDM-cond, UViT-cond, and MultistageEDM-cond), supporting Claim (i). Second, the high $\text{RP}_{between}\text{ Score}$ in Figure 7 (c) provides strong evidence for reproducibility between conditional and unconditional diffusion models, supporting Claim (ii). This can also be observed visually by examining the unconditional generation using Progressivev1 in Figure 7 (a), where images highlighted in the green square share high similarity with conditional counterparts highlighted in the purple square .
Figure 7: Model reproducibility for conditional diffusion model. In this study, we employ conditional diffusion models, specifically U-Net-based (EDM-cond, MultistageEDM-cond) and transformer based (UViT-cond), which we train on the CIFAR-10 dataset using class labels as conditions. Additionally, we select unconditional diffusion models, namely Progressivev1, DDPMv4, and EDMv2, as introduced in Section 2. Figure (a) showcases sample generations from both unconditional and conditional diffusion models (with the "plane" serving as the condition for the latter). Notably, samples within the same row and column originate from the same initial noise. The reproducibility scores between the conditional diffusion models are presented in (b), and between unconditional and conditional diffusion models in (c).
4.2 Model Reproducibility in Solving Inverse Problems
Recently, diffusion models have also demonstrated remarkable results on solving a broad spectrum of inverse problems [12, 11, 51, 9], [^4] Here, the problem is often to reconstruct an unknown signal $\bm{u}$ from the measurements $\bm{z}$ of the form $\bm{z}=\mathcal{A}(\bm{u})+\bm{\eta}$, where $\mathcal{A}$ denotes some (given) sensing operator and $\bm{\eta}$ is the noise. including but not limited to image super-resolution, de-blurring, and inpainting. Motivated by these promising results, we illustrate based upon solving the image inpainting problem using a modified deterministic variant of diffusion posterior sampling (DPS) [11], showcasing that for solving inverse problem using diffusion models:
Model reproducibility largely holds only within the same type of network architectures.
Our claim is supported by the experimental results in Figure 8. Specifically, Figure 8 (a) virtualizes the samples generated from different diffusion models, and Figure 8 (b) presents the similarity matrix of model reproducibility between different models, i.e., U-Net based (DDPMv1, DDPMv2, DDPMv3, DDPMv4, Multistagev1) and Transformer based (DiT, U-ViT) architectures. We note a strong degree of model reproducibility among architectures of the same type (such as U-Net or Transformer), but the model reproducibility score exhibits a notable decrease when comparing any U-Net model to any Transformer-based model.
Figure 8: Model reproducibility for solving inverse problems. In this investigation, we employ various unconditional diffusion models, as introduced in Section 2, which were initially trained on the CIFAR-10 dataset. Our approach involves utilizing a modified deterministic variant of diffusion posterior sampling (DPS), as detailed in Appendix D. Specifically, we focus on the task of image inpainting. Figure (a) presents both the observation $\bm{z}$, unknown signal $\bm{u}$, and generations from different diffusion models. Notably, samples within the same row and column originate from the same initial noise. The reproducibility scores for the generated outputs from different diffusion models under the DPS algorithm are quantitatively analyzed in (b).
4.3 Model Reproducibility in Fine-tuning Diffusion Models
Few-shot image fine-tuning for diffusion models, as discussed in [52, 53, 54, 55], showcases remarkable generalizability. This is often achieved by fine-tuning a small portion of the parameters of a large-scale pre-trained (text-to-image) diffusion model. In this final study, we delve into the impacts of partial model fine-tuning on both model reproducibility and generalizability, by extending our analysis in Section 3. We show that:
Partial fine-tuning reduces reproducibility but improves generalizability in “memorization regime”.
Our claim is supported our results in Figure 9, comparing model fine-tuning and training from scratch of with varying size of the training data, where both models have the same number of parameters. In comparison to training from scratch that we studied in Figure 1(b), fine-tuning specific components of pre-trained diffusion models, particularly the attention layer in the U-Net architecture, yields lower model reproducibility score but higher generalization score in the memorization regime. However, in the generalization regime, partial model fine-tuning has a minor impact on both reproducibility and generalization in the diffusion model. Our result reconfirms the improved generalizability of fine-tuning diffusion models on limited data, but shows a surprising tradeoff in terms of model reproducibility that is worth of further investigations.
Figure 9: Model reproducibility for diffusion model finetuing. In this experiment, we employ DDPMv4. Two distinct training strategies are investigated: "from scratch," denoting direct training on a subset of the CIFAR-10 dataset, and "partial fine-tuning," which involves pretraining on the entire CIFAR-100 dataset [28] followed by fine-tuning only the attention layers of the model on a subset of the CIFAR-10 dataset. The dataset sizes for CIFAR-10 range from $2^{6}$ to $2^{15}$. Importantly, both "from scratch" and "partial fine-tuning" are trained using the same subset of images for each dataset size. Under different dataset sieze, Figure (a) illustrates the reproducibility score between these two strategies and (b) presents the generalization score for them.
5 Discussion and Future Directions
In this work, we conducted an in-depth study of an important but largely overlooked phenomenon in diffusion models, for which we term it as “consistent model reproducibility”. This study raises numerous compelling questions that are worth further exploration, which we discuss in the folllowing
Practical applications.
One such question is the tangible practical benefits of model reproducibility in diffusion models compared to other types of generative models. One major insight gained from this study is that the mapping from the noise space to the image space is a unique one-to-one mapping, regardless of network architectures or training procedures. This observation implies that the noise space contains valuable information that can be utilized for controlled data generative processes using diffusion models. For instance, in text-driven image generation tasks, a deeper understanding of this mapping could facilitate content generation, by leveraging the same text embeddings while varying noise inputs, there are potential applications including adversarial attacks [15] (noise that will always map to adversarial generation), robust defense [16] (noise that robust to adversarial text input), and copyright protection [17, 18] (noise that not memorize data). Moreover, in the context of solving inverse problems, a definitive answer to this question could guide us to select the input noise for reducing the uncertainty and variance in our signal reconstruction  [19, 20, 21].
Theoretical understandings.
Moreover, the strong connection between model reproducibility and generalizability opens an enticing theoretical question for further study. While the phenomenon of “consistent model reproducibility” in the memorization regime has been comprehensively explained in this work, the understanding of this phenomenon in the generalization regime, particularly the identifiable mapping between the noise space and image space, remains an unresolved but valuable question. Recently, the work [48] has developed a framework for analyzing the generalizability of diffusion models, which might offer valuable insights into the reproducibility analysis due to the strong connection between reproducibility and generalizability.
On the other hand, numerous recent studies [56, 57, 58, 59, 60, 61] have explored the connection between diffusion models and the Schrödinger bridge, treating the encoding process from the image space to the noise space as an optimal transport problem. The identifiable relationship could be potentially derived as the solution to the optimal transport problem.
Acknowledgment
HJZ, YFL, and QQ acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, and a gift grant from KLA. LYS and QQ acknowledge support from MICDE Catalyst Grant, and LYS also acknowledges the support from the MIDAS PODS Grant. The authors acknowledge efficient cloud management framework SkyPilot [62] for computing. The authors acknowledge valuable discussions with Prof. Jeffrey Fessler (U. Michigan), Prof. Saiprasad Ravishankar (MSU), Prof. Rongrong Wang (MSU), Prof. Weijie Su (Upenn), Dr. Ruiqi Gao (Google DeepMind), Mr. Bowen Song (U. Michigan), Mr. Xiao Li (U. Michigan), Mr. Zekai Zhang (U. Tsinghua), Dr. Ismail R. Alkhouri (U. Michigan and MSU).
Appendix A Unconditional Diffusion Model
Expanded experiment setting
More detailed settings of the diffusion model we selected are listed in Table 1. With the exception of DiT and UViT, where we implemented and trained them ourselves, all selected diffusion model architectures utilize the author-released models.
Architectural Relationships
For DDPMv1, DDPMv2, and DDPMv7, we adopt the DDPM architecture initially proposed by [1], but we implement it using the codebase provided by [2]. DDPMv3 and DDPMv8, on the other hand, employ DDPM++, an enhanced version of DDPM introduced by [2]. DDPM++ incorporates BigGAN-style upsampling and downsampling techniques, following the work of [63]. DDPMv4, DDPMv5, and DDPMv6 adopt DDPM++(deep), which shares similarities with DDPM++ but boasts a greater number of network parameters. Moving to Multistagev1, Multistagev2, and Multistagev3, these models derive from the Multistage architecture, a variant of the U-Net architecture found in DDPM++(deep). For EDMv1, EDMv2, CT, and CD, the EDM architecture is identical to DDPM++, but they differ in their training parameterizations compared to other DDPM++-based architectures. Finally, UViT and DiT are transformer-based architectures.
Distillation Relationships
CD, Progressivev1, Progressivev2, and Progressivev3 are all diffusion models trained using distillation techniques. CD employs EDM as its teacher model, while Progressivev1, Progressivev2, and Progressivev3 share DDPMv3 as their teacher model. It’s worth noting that these models employ a progressive distillation strategy, with slight variations in their respective teacher models, as elaborated in [38].
Initial Noise Consistency
However, it is important to note a nuanced difference related to the noise perturbation kernels. Specifically, for VP and subVP noise perturbation kernels, we define the noise space as $\mathcal{E}=\mathcal{N}(\bm{0},\bm{I})$, whereas the VE noise perturbation kernel introduces a distinct noise space with $\mathcal{E}=\mathcal{N}(\bm{0},\bm{\sigma}_{\text{max}}\cdot I)$, where $\sigma_{\text{max}}$ is predefined. So during the experiment, we sample 10K initial noise $\bm{\epsilon}_{\text{vp, subvp}}\sim\mathcal{N}(\bm{0},\bm{I})$ for the sample generation of diffusion models with VP and subVP noise perturbation kernel. For diffusion models with VE noise perturbation kernel, the initial noise is scaled as $\bm{\epsilon}_{\text{ve}}=\sigma_{\text{max}}\bm{\epsilon}_{\text{vp, subvp}}$.
Additionally, it’s worth mentioning that for all 8x8 image grids shown in the Figure 1, 12, 13, 14, 15, 16, 18, 19, 21 no matter for the unconditional diffusion model, conditional diffusion model, diffusion model for the inverse problem, or fine-tuning diffusion model, we consistently employ the same 8x8 initial noise configuration. The same setting applies to 10k initial noises for reproducibility score. This specific design is for more consistent results between different variants of diffusion models (e.g., we could clearly find the relationship between the unconditional diffusion model and conditional diffusion model by comparing Figure 12 and Figure 15, 16).
Further discussion
In Figure 12, we provide additional visualizations, offering a more comprehensive perspective on our findings. For a deeper understanding of our results, we present extensive quantitative data in Figure 11 and Figure 10. Building upon the conclusions drawn in Section 2, we delve into the consistency of model reproducibility across discrete and continuous timestep settings. To illustrate, we compare DDPMv1 and DDPMv2, demonstrating that model reproducibility remains steadfast across these variations.Moreover, it’s worth noting that while all reproducibility scores surpass a threshold of 0.6, signifying robust model reproducibility, some scores do exhibit variations. As highlighted in Figure 10, we observe that similar architectures yield higher reproducibility scores (e.g., DDPMv1-8), models distilled from analogous teacher models exhibit enhanced reproducibility (e.g., Progressivev1-3), and models differing solely in their ODE samplers also display elevated reproducibility scores (e.g., DDPMv4, DDPMv5).We hypothesize that the disparities in reproducibility scores are primarily attributed to biases in parameter estimation. These biases may arise from factors such as differences in architecture, optimization strategies, and other variables affecting model training.
Table 1: Comprehensive unconditional reproducibility experiment settings
Name Architecture SDE Sampler Continuous Distillation
DDPMv1 DDPM VP DPM-Solver
DDPMv2 DDPM VP DPM-Solver
DDPMv3 DDPM++ VP DPM-Solver
DDPMv4 DDPM++(deep) VP DPM-Solver
DDPMv5 DDPM++(deep) VP ODE
DDPMv6 DDPM++(deep) sub-VP ODE
DDPMv7 DDPM sub-VP ODE
DDPMv8 DDPM++ sub-VP ODE
Multistagev1 Multistage (3 stages) VP DPM-Solver
Multistagev2 Multistage (4 stages) VP DPM-Solver
Multistagev3 Multistage (5 stages) VP DPM-Solver
EDMv1 EDM VP Heun-Solver
EDMv2 EDM VE Heun-Solver
UViT UViT VP DPM-Solver
DiT DiT VP DPM-Solver
CD EDM VE 1-step
CT EDM VE 1-step
Progressivev1 DDPM++ VP DDIM (1-step)
Progressivev2 DDPM++ VP DDIM (16-step)
Progressivev3 DDPM++ VP DDIM (64-step)
Figure 10: Comprehensive reproducibility score among different unconditional diffusion model settings.
Figure 11: Comprehensive MAE score among different unconditional diffusion model settings.
(a) DDPMv1
(b) DDPMv2
(c) DDPMv3
(d) DDPMv5
(e) DDPMv6
(f) DDPMv7
(g) DDPMv8
(h) EDMv1
(i) EDMv2
(j) DiT
(k) CD
(l) Progressivev1
(m) Progressivev2
(n) Progressivev3
Figure 12: Comprehensive samples visulization for unconditional diffusion model
Appendix B Theoretical Analysis
This section mainly focus on the proof of Theorem 1 in Section 3. The proof is mainly built upon recent works in [32].
Proof.
As the background, let $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$ be the perturbation kernel of diffusion model, which is a continuous process gradually adding noise from original image $\bm{x}_{0}$ to $\bm{x}_{t}$ along the timestep $t\in[0,1]$. Both $s_{t}=s(t),\sigma_{t}=\sigma(t)$ here are simplified as scalar functions of $t$ to control the perturbation kernel. It has been shown that this perturbation kernel is equivalent to a stochastic differential equation $\text{d}\bm{x}=f(t)\bm{x}\text{d}t+g(t)\text{d}\bm{\omega}_{t}$, where $f(t),g(t)$ are a scalar function of $t$. The relations of $f(t),g(t)$ and $s_{t},\sigma_{t}$ are:
$s_{t}=\text{exp}(\int_{0}^{t}f(\xi)\text{d}\xi),\ \ \text{and}\ \sigma_{t}=\sqrt{\int_{0}^{t}\frac{g^{2}(\xi)}{s^{2}(\xi)}\text{d}\xi}$ (3)
Given a dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$ with $N$ images, we model the original dataset distribution $p_{\text{data}}$ as multi-Dirac distribution, $p_{\text{data}}(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}-\bm{y}_{i})$, the distribution of perturbed image $\bm{x}$ at random timestep $t$ could be calculated as:
$p_{t}(\bm{x})$ $=\int_{\mathbb{R}^{d}}p_{t}(\bm{x}|\bm{x}_{0})p_{\text{data}}(\bm{x}_{0})\text{d}\bm{x}_{0}$ (4)
$=\int_{\mathbb{R}^{d}}p_{\text{data}}(\bm{x}_{0})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (5)
$=\int_{\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}_{0}-\bm{y}_{i})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (6)
$=\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\delta(\bm{x}_{0}-\bm{y}_{i})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (7)
$=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$ (8)
Let us consider the noise prediction loss used generally across various diffusion model works:
$\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$ $=\mathbb{E}_{\bm{x}\sim p_{t}(\bm{x})}[|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)||^{2}]$ (9)
$=\int_{\mathbb{R}_{d}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})||\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)||^{2}\text{d}\bm{x}$ (10)
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\textbf{I})$ is defined follow the perturbation kernel $p_{t}(\bm{x}|\bm{x}_{0})=\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$:
$\bm{x}=s_{t}\bm{y}_{i}+s_{t}\sigma_{t}\bm{\epsilon}\Rightarrow\bm{\epsilon}=\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}$ (11)
And $\bm{\epsilon}_{\bm{\theta}}$ is a "denoiser" network for learning the noise $\bm{\epsilon}$.
So plugging Eq. 11 into 10, we could reparameterization the loss as:
$\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)=\int_{\mathbb{R}_{d}}\underbrace{\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})||\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)-\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}||^{2}}_{=:\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)}\text{d}\bm{x}$ (12)
Eq. 12 means we could minimize $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$ by minimizing $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ for each $\bm{x}$. And to find the "optimal denoiser" $\bm{\epsilon}^{*}_{\bm{\theta}}$ that minimize the $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ for every given $\bm{x},t$:
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\text{arg}\ \text{min}_{\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ (13)
This is a convex optimization problem; the solution could be solved by setting the gradient of $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ w.r.t $\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)$ to zero:
$\nabla_{\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)}[\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)]=0$ (14)
$\Rightarrow$ $\nabla_{\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)}[\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})||\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)-\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}||^{2}]=0$ (15)
$\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})[\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)-\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}]=0$ (16)
$\Rightarrow$ $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}]$ (17)
It is obvious that the optimal denoiser $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ is a function only depend on the perturbation kernel parameter $s_{t},\sigma_{t}$ and the dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$.
Given the assumption that the denoiser $\bm{\epsilon}_{\bm{\theta}}$ could converge to the optimal denoiser $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ we use a deterministic ODE sampler to generate images, then mapping $f:\mathcal{E}\mapsto\mathcal{I}$, from gaussion noise space $\mathcal{E}$ to image space $\mathcal{I}$ is an invertible mapping and the inverse mapping $f^{-1}$ is a unique identifiable encoding.
The mapping $f$ is only determined by the $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)$ and the ODE sampler. Take the probability flow ODE sampler [25] as an example, the ODE is given as:
$\frac{\text{d}\bm{x}_{t}}{\text{d}t}=f(t)\bm{x}_{t}+\frac{g^{2}(t)}{2s_{t}\sigma_{t}}\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t};t)$ (18)
The mapping $f$ given the optimal denoiser $\bm{\epsilon}^{*}_{\bm{\theta}}$ could be determined as:
$f:\mathcal{E}\mapsto\mathcal{I}:=\left\{\begin{aligned} &\frac{\text{d}\bm{x}_{t}}{\text{d}t}=f(t)\bm{x}_{t}+\frac{g^{2}(t)}{2s_{t}\sigma_{t}}\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x}_{t};t)\text{d}t\\ &\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}]\\ &\bm{x}_{t=0}\in\mathcal{I},\ \bm{x}_{t=1}\in\mathcal{E},\ t:1\mapsto 0\\ \end{aligned}\right.$ (19)
The mapping $f$ is invertible and the inverse mapping $f^{-1}$ is defined as:
$f^{-1}:\mathcal{I}\mapsto\mathcal{E}:=\left\{\begin{aligned} &\frac{\text{d}\bm{x}_{t}}{\text{d}t}=f(t)\bm{x}_{t}+\frac{g^{2}(t)}{2s_{t}\sigma_{t}}\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x}_{t};t)\text{d}t\\ &\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}]\\ &\bm{x}_{t=0}\in\mathcal{I},\ \bm{x}_{t=1}\in\mathcal{E},\ t:0\mapsto 1\\ \end{aligned}\right.$ (20)
For the mapping $f^{-1}$ a clear image $\bm{x}_{t=0}$ would generate a deterministic embedding noise $\bm{x}_{t=1}$. If ignoring the discretization error from the ODE sampler, the reverse mapping $f$ could numerically also start from $\bm{x}_{t=1}$ to generate the specific image $\bm{x}_{t=0}$ [4]. So the encoding is unique. All the above processes could be expressed by the optimal denoiser and ODE solver, which is identifiable. In conclusion, the mapping $f$ of the diffusion model is uniquely identifiable encoding theoretically. ∎
Figure 13: Visualization between theoretical and experimental results.
Figure 14: Visualization between theoretical and experimental results.
Extended Experiment Setting
Similar to what is illustrated in Figure 5, we employ the theoretical generation process, facilitated by the inverse mapping $f^{-1}$. To expedite the sampling speed, we leverage the DPM-Solver [40].For a more comprehensive view of our results, we present additional visualizations in Figure 13 and Figure 14. In these experiments, we train UNet models with varying numbers of channels on subsets of the CIFAR-10 dataset, each comprising different training samples. Our standard batch size for all experiments is set at 128, and we continue training until the generated samples reach visual convergence, characterized by minimal changes in both appearance and semantic information.
Appendix C Conditional Diffusion Model
Extended Experiment setting To investigate the reproducibility of the conditional diffusion model, we opted for three distinct architectures: the conditional EDM [32], conditional multistage EDM [35], and conditional U-ViT [27]. Our training data consisted of the CIFAR-10 dataset, with the class labels serving as conditions. It’s worth noting that the primary distinction between EDM and multistage EDM lies in the architecture of the score function. Conversely, the contrast between EDM and conditional U-ViT extends beyond architectural differences to encompass conditional embeddings. Specifically, EDM transforms class labels into one-hot vectors, subjects them to a single-layer Multilayer Perceptron (MLP), and integrates the output with timestep embeddings. In contrast, U-ViT handles class labels by embedding them through a trainable lookup table, concatenating them with other inputs, including timestep information and noisy image patches represented as tokens. For all three architectures, we pursued training until convergence was achieved, marked by the lowest FID. The DPM-Solver was employed for sampling purposes. To generate samples, we employed the same 10K initial noise distribution as utilized in the unconditional setting (refer to Section 2.1). For each such initial noise instance, we generated 10 images, guided by 10 distinct classes, resulting in a total of 100K images.
Discussion The observed reproducibility between the unconditional diffusion model and the conditional diffusion model presents an intriguing phenomenon. It appears that the conditional diffusion model learns a mapping function, denoted as $f_{c\in\mathcal{C}}:\mathcal{E}\mapsto\mathcal{I}_{c\in\mathcal{C}}$, which maps from the same noise space $\mathcal{E}$ to each individual image manifold $\mathcal{I}_{c\in\mathcal{C}}$ corresponding to each class $c$. In contrast, the mapping of the unconditional diffusion model, denoted as $f:\mathcal{E}\mapsto\mathcal{I}$, maps the noise space to a broader image manifold $\mathcal{I}\subset\bigcup_{c\in\mathcal{C}}\mathcal{I}_{c}$. A theoretical analysis of this unique reproducibility relationship holds the promise of providing valuable insights.
Currently, our research is exclusively focused on the conditional diffusion model. It raises the question of how the reproducibility phenomenon manifests in the context of the text-to-image diffusion model [3, 7, 8], where the conditioning factor is not confined to finite classes but instead involves complex text embeddings.
As illustrated in Figure 15 and Figure 16, our previous comparisons were made with the same initial noise and class conditions. However, when comparing the same model with identical initial noise but different class conditions, we uncovered intriguing findings. For instance, the first row and column images in Figure 15 (i) and (l) exhibited remarkable similarity in low-level structural attributes, such as color, despite differing in semantics. This observation is consistent with findings in Figure 21, where we explored generation using diffusion models trained on mutually exclusive CIFAR-100 and CIFAR-10 datasets. These findings bear a striking resemblance to the conclusions drawn in [61], which also demonstrated a similar phenomenon in a simplified scenario, where $\mathcal{I}$ follows a Gaussian distribution. To gain a deeper understanding of reproducibility and the phenomena mentioned in this paragraph, leveraging optimal transport methods (e.g., Schrödinger bridge [56, 57, 58, 59, 60]) holds significant potential.
(a) EDM Class0
(b) MultistageEDM Class0
(c) U-ViT Class0
(d) EDM Class1
(e) MultistageEDM Class1
(f) U-ViT Class1
(g) EDM Class2
(h) MultistageEDM Class2
(i) U-ViT Class2
(j) EDM Class3
(k) MultistageEDM Class3
(l) U-ViT Class3
(m) EDM Class4
(n) MultistageEDM Class4
(o) U-ViT Class4
Figure 15: Visualization of conditional diffusion model generations (class 0 - 4).
(a) EDM Class5
(b) MultistageEDM Class5
(c) U-ViT Class5
(d) EDM Class6
(e) MultistageEDM Class6
(f) U-ViT Class6
(g) EDM Class7
(h) MultistageEDM Class7
(i) U-ViT Class7
(j) EDM Class8
(k) MultistageEDM Class8
(l) U-ViT Class8
(m) EDM Class9
(n) MultistageEDM Class9
(o) U-ViT Class9
Figure 16: Visualization of conditional diffusion model generations (class 5 - 9).
Appendix D Diffusion Model for Solving Inverse Problem
To explore the reproducibility of diffusion models in solving inverse problems, we adopted the Diffusion Posterior Sampling (DPS) strategy proposed by Chung et al. [11]. Our adaptation involved a slight modification of their algorithm, specifically by eliminating all sources of stochasticity within it. Additionally, we employed the DPM-Solver for Diffusion Posterior Sampling.
Extended Experiment setting To explore the reproducibility of diffusion models in solving inverse problems, we adopted the Diffusion Posterior Sampling (DPS) strategy proposed by Chung et al. [11]. Our adaptation involved a slight modification of their algorithm, specifically by eliminating all sources of stochasticity within it. Additionally, we employed the DPM-Solver for Diffusion Posterior Sampling: Algorithm 1, with $N_{\text{dps}}=34$ posterior samping steps, 33 iterations for 3rd order DPM-Solver, 1 for 1st order DPM-Solver, thus 100 function evaluations. We also set all $\xi_{i}=1.$
For the task involving image inpainting on the CIFAR-10 dataset, we applied two square masks to the center of the images. One mask measured 16 by 16 pixels, covering 25% of the image area, and the other measured 25 by 25 pixels, covering 61% of the image area. We denoted these as "easy inpainting" and "hard inpainting" tasks. In Figure 8 and Figure 17, we utilized the "easy inpainting" scenario with a specific observation $\bm{z}$ as illustrated in the figure. In Figure 20, we considered both the "easy inpainting" and "hard inpainting" tasks. We also employed 10K distinct initial noise and their corresponding 10K distinct observations $\bm{z}$ to calculate the reproducibility score, as presented in Figure 20. Additional visualizations for Figure 8 and Figure 17 are provided in Figure 18 and Figure 19.
Algorithm 1 Determinsitic DPS with DPM-Solver.
1:$N_{\text{dps}}$, $\bm{u}$,$f(t)$,$g(t)$, $s_{t}$, $\sigma_{t}$, $\xi_{i}\}_{i=1}^{N_{\text{dps}}}$
2:${\bm{x}}_{N_{\text{dps}}}\sim{\mathcal{N}}(\bm{0},\bm{I})$
3:for $i=N_{\text{dps}}$ to $q$ do
4:     $\hat{\bm{x}}_{0}=\dfrac{1}{f(i)}\left(\bm{x}_{i}-\dfrac{g^{2}(i)}{s_{i}\sigma_{i}}\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{i},i\right)\right)$
5:     ${\bm{x}}^{\prime}_{i-1}\leftarrow\text{Dpm-Solver}({\bm{x}}_{i},i)$
6:     $\bm{x}_{i-1}\leftarrow\bm{x}^{\prime}_{i-1}-\xi_{i}\nabla_{\bm{x}_{i}}||\bm{u}-\mathcal{A}\left(\hat{\bm{x}}_{0}\right)||^{2}_{2}$
7:end for
8:return $\hat{\mathbf{x}}_{0}$
Discussion Reproducibility is a highly desirable property when employing diffusion models to address inverse problems, particularly in contexts such as medical imaging where it ensures the reliability of generated results. As observed in Figure 17, the reproducibility scores vary for different observations $\bm{z}$, and the decrease in reproducibility differs across various architecture categories. For instance, when considering observation $\bm{z}_{1}$, the reproducibility scores across different architecture categories remain above 0.5, whereas for $\bm{z}_{3}$, they fall below 0.3. Since the choice of observation $\bm{z}$ also significantly impacts reproducibility, we conducted a complementary experiment presented in Figure 20. In this experiment, for each initial noise instance, we employed a different observation $\bm{z}$. From the results, it is evident that reproducibility decreases between different categories of diffusion models. Furthermore, reproducibility diminishes as the inpainting task becomes more challenging, with "hard inpainting" being more demanding than "easy inpainting."
Here is an intuitive hypothesis of the decreasing reproducibility:
The update step of Diffusion Posterior Sampling (DPS), is constrained by the data consistency through the following equation:
$\bm{x}_{i-1}\leftarrow\bm{x}^{\prime}_{i-1}-\xi_{i}\nabla_{\bm{x}_{i}}||\bm{u}-\mathcal{A}\left(\hat{\bm{x}}_{0}\right)||^{2}_{2}$ (21)
Where $\hat{\bm{x}}_{0}=\dfrac{1}{f(i)}\left(\bm{x}_{i}-\dfrac{g^{2}(i)}{s_{i}\sigma_{i}}\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{i},i\right)\right)$, we could show that:
$\xi_{i}\nabla_{\bm{x}_{i}}||\bm{z}-\mathcal{A}\left(\hat{\bm{x}}_{0}\right)||^{2}_{2}$ $=\frac{\partial\mathcal{A}\left(\hat{\bm{x}}_{0}\right)}{\partial\bm{x}_{i}}\left(\mathcal{A}\left(\hat{\bm{x}}_{0}\right)-\bm{z}\right)$ (22)
$=\frac{\partial\mathcal{A}\left(\hat{\bm{x}}_{0}\right)}{\partial\hat{\bm{x}}_{0}}\dfrac{\partial\hat{\bm{x}}_{0}}{\partial\bm{x}_{i}}\left(\mathcal{A}\left(\hat{\bm{x}}_{0}\right)-\bm{z}\right)$ (23)
$=\dfrac{1}{f(i)}\frac{\partial\mathcal{A}\left(\hat{\bm{x}}_{0}\right)}{\partial\hat{\bm{x}}_{0}}\left(1-\dfrac{g^{2}(i)}{s_{i}\sigma_{i}}\dfrac{\partial\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{i},i\right)}{\partial x_{i}}\right)\left(\mathcal{A}\left(\hat{\bm{x}}_{0}\right)-\bm{z}\right)$ (24)
This analysis highlights that the unconditional diffusion model is reproducible as long as the function $\bm{\epsilon}_{\bm{\theta}}$ is reproducible. However, for the diffusion model used in inverse problems to be reproducible, both the function $\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{t},t\right)$ and its first-order derivative with respect to $\bm{x}_{t}$ must be reproducible. In other words, the denoiser should exhibit reproducibility not only in its results but also in its gradients. Combining the findings in Figure 20, we can infer that for similar architectures, reproducibility also extends to the gradient space $\dfrac{\partial\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{t},t\right)}{\partial\bm{x}_{t}},$ which may not hold true for dissimilar architectures. Ensuring reproducibility in the gradient space should thus be a significant focus for achieving reproducibility in diffusion models for solving inverse problems.
Additionally, it’s worth noting that the data $\bm{x}_{t}$ passed into the denoiser $\bm{\epsilon}_{\bm{\theta}}\left(\bm{x}_{t},t\right)$ is always out-of-distribution (OOD) data, especially in tasks like image inpainting. Consequently, the reproducibility of OOD data $\bm{x}_{t}$ is also crucial for achieving reproducibility in diffusion models for solving inverse problems.
(a) observation $\bm{z}_{1}$
(b) observation $\bm{z}_{2}$
(c) observation $\bm{z}_{3}$
Figure 17: Visualization of inverse problem solving with different observations
Figure 18: More visualization results for Figure 8 and Figure 17
Figure 19: More visualization results for Figure 8 and Figure 17
Figure 20: Extended experiments on image impainting for reproducibility score.
Appendix E Fine-tuning Diffusion Model
Extended Experiment setting In our investigation of reproducibility during fine-tuning, we first trained an unconditional diffusion model using EDM [32] on the CIFAR-100 dataset [28]. All the fine-tuned models discussed in this section were pre-trained on this model. Subsequently, we examined the impact of dataset size by conducting fine-tuning on the EDM using varying numbers of CIFAR-10 images: 64, 1024, 4096, 16384, and 50000, respectively. Building upon the findings in [54], which indicate that fine-tuning the attention blocks is less susceptible to overfitting, we opted to target all attention layers for fine-tuning in our experiments. For comparison purposes, we also trained a diffusion model from scratch on the CIFAR-10 dataset, using the same subset of images. All models were trained for the same number of training iterations and were ensured to reach convergence, as evidenced by achieving a low Fréchet Inception Distance (FID) and maintaining consistent mappings from generated samples. The training utilized a batch size of 128 and did not involve any data augmentation.
Extended Results Additional generations produced by both the "from scratch" diffusion models and the fine-tuned diffusion models are presented in Figure 21, encompassing various training dataset sizes. A notable observation arises when comparing the fine-tuned diffusion model’s generation using 4096 and 50000 data samples. Even with this limited dataset, the fine-tuned diffusion model demonstrates a remarkable ability to approximate the target distribution. This suggests that the fixed portion of the diffusion model, containing information from the pre-trained CIFAR-100 dataset, aids the model in converging to the target distribution with less training data. In contrast, when attempting to train the diffusion model from scratch on CIFAR-10, even with 16384 data samples, it fails to converge to the target distribution. Additionally, despite the distinct nature of CIFAR-100 and CIFAR-10, their generations from the same initial noise exhibit striking similarities (Figure 21). This similarity might be a contributing factor explaining how the pre-trained CIFAR-100 diffusion model assists in fine-tuning the diffusion model to converge onto the CIFAR-10 manifold with reduced training data.
Figure 21: More visualization of finetuning diffusion models